OpenStudy (anonymous):

What is the volume of the prism? _____cm3

OpenStudy (mathstudent55):

\(\LARGE V_{prism} = A_{base} \times h\)

OpenStudy (anonymous):

\[648cm ^{3}?\]

OpenStudy (anonymous):

actual answer is really \[324cm^{3}\]

OpenStudy (mathstudent55):

In this prism, the base is the triangular face on the left side. The base is a right triangle with base = 6 cm and height = 12 cm. Then the prism itself has a height of 9 cm. The volume of the prism is the area of the base multiplied by the height of the prism. \(\LARGE V = \color{red}{A_{triangle}}\color{green}{h_{prism}} \) \(\LARGE V = \color{red}{\dfrac{1}{2}b_{triangle}h_{triangle}}\color{green}{h_{prism}} \) \(\LARGE V = \color{red}{\dfrac{1}{2}(6~cm)(12~cm)}\color{green}{(9~cm)}\) \(\LARGE V = \color{red}3 \color{green}2\color{red}4 ~\color{green}c\color{red}m^\color{green}3\)